A metal tape correct at 30 degrees is used to measure the length of field on a day when the temperature is 20 degrees. How does this affect the result.? What is the error measuring a distance of 10meters? The linear expansivity of the metal is 18x10 raise to power of - 6 per k.
1
Expert's answer
2019-03-20T11:48:22-0400

The linear thermal contraction of metal tape

"\Delta l=l_0\alpha\Delta T""=(1\:\rm{m})\times(18\times10^{-6}\:\rm{K^{-1}})\times (10\:\rm{K})=18\times10^{-5}\:\rm{m}"

So, the error measuring a distance of 10 m is

"\rm{error}=(10\:\rm{m})\times(18\times10^{-5}\:\rm{m})=18\times10^{-4}\:\rm{m}=1.8\:\rm{mm}"